The motion of a rocket burning its fuel at a constant rate while moving through empty
interstellar space can be described by
x(t) = uext + uex
1
b
− t

ln(1 − bt).
uex and b are constants (uex is the exhaust velocity of the gases at the tail of the rocket,
and b is proportional to the rate of fuel consumption). For this question, you should
know that ln is the inverse of the exponential function, that is, ln(e
x
) = x. Also,
d
dx ln x =
1
x
. [10]
a) What happens when t > 1/b? Explain.
b) Find a formula for the instantaneous velocity of the rocket.
c) Find a formula for the instantaneous acceleration.
d) Suppose that a rocket with uex = 3.0 × 103 m/s and b = 7.5 × 10−3
s
−1
takes 120
s to burn all its fuel. What is the instantaneous velocity at t = 0 s? At t = 120 s?
1
Expert's answer
2020-10-07T07:31:29-0400

As per the given question,

"x(t)=u_{ex}\left(\dfrac{1}{b}-t\right)\ln\left(1-bt\right)+u_{ex}t"

a) As per the question, if "t>\frac{1}{b}"

"x(t)=u_{ex}\left(\dfrac{1}{b}-t\right)\ln\left(1-bt\right)+u_{ex}t"

so tb>1 hence first term will not defined because ln is not defined for negative value.


b)

Now taking the differentiation with respect to t

"\frac{x(t)}{dt}=\frac{d }{dx}[(u_{ex}\left(\frac{1}{b}-t\right)\ln\left(1-bt\right)+u_{ex}t)]"


"\Rightarrow" "dv=u_{ex}\frac{d}{dt}[(\left(\frac{1}{b}-t\right)\ln\left(1-bt\right)]+u_{ex}"


"\Rightarrow" "dv=u_{ex}[\frac{d}{dt}[(\frac{1}{b}-t)\ln\left(1-bt\right)+[(\frac{1}{b}-t)\frac{d}{dt}\ln(1-bt)]+u_{ex}"


"\Rightarrow dv= u_{ex}\left(-\ln\left(1-bt\right)-\dfrac{b\left(\frac{1}{b}-t\right)}{1-bt}\right)+t"

"\Rightarrow dv=-u_{ex}\ln\left(1-bt\right)-\dfrac{bu_{ex}\left(\frac{1}{b}-t\right)}{1-bt}+u"

"\Rightarrow dv= -u\ln\left(1-bt\right)"

c) For the instantaneous acceleration,

"da=\frac{dv}{dt}"


"\Rightarrow da=\frac{d}{dt} \left( -u_{ex} \ln \left(1-bt\right) \right)"


"\Rightarrow da=-u_{ex}\frac{d}{dt}\left( \ln(1-bt)\right)"


"\Rightarrow da=-u_{ex}(\frac{1}{1-bt})\frac{d}{dt}(1-bt)"


"\Rightarrow da=\dfrac{bu_{ex}}{1-bt}"


d) Now it is given

"u_{ex} = 3.0 \u00d7 10^3 m\/s"


"b = 7.5 \u00d7 10^{\u22123}\/sec"

"t=0"

"t=120 sec"

"\Rightarrow dv= -u\ln\left(1-bt\right)"

at t=0

"dv=3.0\times 10^3\ln(1-7.5\times 10^{-3}\times 0)"

"\Rightarrow dv=3.0\times 10^3\ln(1)"

"\Rightarrow dv=0"

Now,

at t=120

"dv=3.0\times 10^3\ln(1-7.5\times 10^{-3}\times 120)"

"\Rightarrow dv=3.0\times 10^3\ln(1-0.016)"

"\Rightarrow dv=0.016\times 3\times 10^3 m/sec"

"\Rightarrow dv=-48m/sec"